# Uncertainty Quantification for Retrieval-Augmented Reasoning

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Figure 1. $\text{R}^{2}$C overview. Given a user query, the agent (LLM) first generates the most-likely reasoning path leading to the most-likely response (left, yellow). To estimate uncertainty, $\text{R}^{2}$C creates multiple perturbed generations by randomly altering states in the reasoning path (middle, gray). The uncertainty score is then derived via majority voting. $\text{R}^{2}$C significantly outperforms established UQ methods and achieves significant improvements on two downstream tasks: abstention and model selection.
Retrieval-augmented generation (RAG) is widely used for knowledge-intensive tasks, but remains limited in addressing complex multi-step reasoning (Jin et al., 2025; Soudani et al., 2024c; Salemi and Zamani, 2024; Soudani et al., 2024b). Recent work has explored combining RAG with reasoning, where LLMs are prompted or trained to use search engines as tools during their reasoning process; a paradigm referred to as retrieval-augmented reasoning (RAR) (Sun et al., 2025; Tran et al., 2025; Islam et al., 2024; Jin et al., 2025). However, RAR models are still prone to producing incorrect responses, due to issues such as retrieving irrelevant documents in early steps, misinterpreting retrieved content, or misusing internal knowledge. Therefore, ensuring the trustworthiness of RAR outputs has become a critical challenge.
Uncertainty quantification (UQ) is a widely studied task in machine learning, aimed at assessing the reliability of model outputs by measuring the degree of uncertainty (or lack of confidence) a model has in its predictions (Malinin and Gales, 2021; Kadavath et al., 2022; Kuhn et al., 2023; Hou et al., 2024; Ye et al., 2024; Harrison et al., 2024). Recent methods of estimating the uncertainty of LLM outputs are designed for settings where the input consists solely of a query, meaning the LLM itself is the only source of uncertainty (Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024). The limited work on UQ for RAG (Soudani et al., 2025; Perez-Beltrachini and Lapata, 2025) incorporates the document–response relationship into the uncertainty score, but these methods are only applicable for simple RAG settings, where documents are retrieved once and inserted into the input prompt for generation. As a result, existing UQ approaches are suboptimal for RAR.
@@
In this paper, we propose Retrieval-Augmented Reasoning Consistency ($\text{R}^{2}$C), a novel UQ method that, unlike previous approaches, accounts for multiple sources of uncertainty in RAR. The central idea of $\text{R}^{2}$C is to allow the model to explore diverse reasoning paths, queries, and documents and then measure the consistency of the resulting final answers. This is achieved by modeling RAR as a Markov Decision Process (MDP) and perturbing this process in a controlled way through a set of perturbation actions across various states. Three perturbation actions are designed to influence query generation, document retrieval, and LLMs’ thinking process. These perturbations enable models to arrive at diverse final responses for uncertain generations. The uncertainty score is then obtained by measuring consistency of the generated answers using majority voting; see Figure 1.
@@
We conduct our experiments across multiple datasets and five RAR models. Our experiments show that $\text{R}^{2}$C significantly outperforms existing LLM-specific UQ methods, achieving on average more than a 5% improvement in AUROC compared to the state-of-the-art UQ methods. We further extrinsically evaluate $\text{R}^{2}$C on two downstream tasks: (i) Abstention: [^1] Also referred to as selective prediction in the literature (Duan et al., 2025; Xin et al., 2021). the task of generating ‘I don’t know’ when the model is uncertain about its output (Feng et al., 2024; Madhusudhan et al., 2025), and (ii) Model Selection: [^2] Also referred to as selection-based model ensemble in the literature (Jiang et al., 2023a). the task of selecting a final answer from a pool of candidates generated by multiple systems (Chen et al., 2025b; Guha et al., 2024). Our experimental results indicate that $\text{R}^{2}$C delivers statistically significant gains over existing approaches: in Abstention, it achieves roughly 5% improvements on both F1Abstain and AccAbstain; in Model Selection, it increases exact match by about 7% relative to single RAR models and by about 3% compared to selection model baselines.
@@
Given the strong performance of $\text{R}^{2}$C in both direct evaluation and extrinsic evaluation on downstream tasks, we investigate the factors that contributes to its effectiveness. We show that $\text{R}^{2}$C retrieves on average 25 unique documents for each score, compared to 16 documents retrieved by other UQ methods. It also achieves a query diversity of $0.35$ compared to $0.30$ other methods, measured by the inverse of the average pairwise cosine similarity between queries (Cox et al., 2021; Zhang et al., 2025). This diversity in queries and documents demonstrates that $\text{R}^{2}$C generates diverse, yet relevant reasoning paths through our controlled perturbation mechanism. As a result, this enables the method to achieve uncertainty scores comparable to baseline approaches while requiring only about 3 generations on average, 2.5 times fewer token generations than the 10 used by the baselines. This highlights that $\text{R}^{2}$C is not only the most effective method of its kind but also a relatively more efficient uncertainty quantification approach. To summarize, the main contributions of this paper are:
@@ 2. Related Work
Retrieval-Augmented (Reasoning) Models. RAG is a framework that combines the strengths of retrieval models and generative models (Cuconasu et al., 2024). Broadly, RAG can be implemented in different ways. In the retrieve-then-generate paradigm, relevant documents are first retrieved based on the user’s input and then incorporated into the model’s prompt (Mallen et al., 2023; Soudani et al., 2024a). In contrast, Active RAG allows retrieval to occur throughout the generation process, either in fixed intervals or dynamically, whenever additional information is needed (Trivedi et al., 2023; Jiang et al., 2023b; Su et al., 2024). Retrieval-Augmented Reasoning (RAR) is a recent extension of RAG that integrates retrieval with reasoning, aiming to improve the interaction between LLMs and retrievers (Sun et al., 2025; Tran et al., 2025). For example, SelfAsk (Press et al., 2023) decomposes complex questions into follow-up queries and intermediate answers, while ReAct (Yao et al., 2023) defines a set of actions, such as search, look up, and finish, to structure interactions with external resources. More recent and effective models such as ReSearch (Chen et al., 2025c) and Search-R1 (Jin et al., 2025) are explicitly trained to seamlessly integrate external resources into reasoning. Despite their effectiveness on complex queries, RAR models remain prone to errors, such as retrieving irrelevant documents, misinterpreting content, or misusing internal knowledge.
@@
Uncertainty Quantification. Broadly, existing UQ methods can be divided into two categories: white-box approaches, which leverage token-level probabilities and entropy (Kadavath et al., 2022; Kuhn et al., 2023; Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024; Soudani, 2025), and black-box approaches, which rely only on the final textual outputs (Lin et al., 2024; Tian et al., 2023). Most UQ methods focus on question answering and view the LLM as the only source of uncertainty, but in RAG, the retriever also contributes its own uncertainty, making this assumption incomplete. Limited research has explored UQ for RAG by modeling the document-response link, either via axioms (Soudani et al., 2025) or utility models (Perez-Beltrachini and Lapata, 2025). Some path-based approaches focus on assessing the consistency of reasoning paths in reasoning tasks (Wang et al., 2023; Qi et al., 2025; Li et al., 2025b). However, these methods do not extend naturally to RAR that involves repeated retrieval during reasoning. Recent work has studied uncertainty propagation in multi-step decision-making by combining uncertainties from intermediate steps.SAUP (Zhao et al., 2025) learns aggregation weights to merge per-step uncertainties, but it relies on ground-truth labels from the test domain. In contrast, we propose $\text{R}^{2}$C, a method that accumulates uncertainty over the entire reasoning path, while considering different sources of uncertainty including the retriever and the generator.
@@ 3. Preliminaries
Consistency-based Uncertainty Quantification. The core idea of consistency-based methods is to generate multiple responses for a given query by varying either the input prompt or the temperature parameter used in stochastic decoding (Korikov et al., 2025). The pairwise similarity among these responses is then computed and aggregated into a single uncertainty score (Lin et al., 2024; Bakman et al., 2025). Formally, consider a model $\pi$ parameterized with $\theta$, generates the most-likely response $r$ by setting the sampling temperature to less than one. Then, $B$ additional responses $R_{G}=\{r^{b}\}_{b=1}^{B}$ are sampled using various sampling strategies such as increasing the temperature (Wang et al., 2023), changing the input (Korikov et al., 2025), or altering the reasoning path (Qi et al., 2025; Li et al., 2025b). A transformation function $\mathcal{\phi}$ is then applied to convert $R_{G}$ into an uncertainty score. Our method builds on the reasoning path perturbation approach and employs Majority Voting (Wang et al., 2023) as the transformation function, where the degree of consistency is measured by the proportion of sampled responses that match the most-likely response:
@@ 4. Methodology
We propose Retrieval-Augmented Reasoning Consistency, $\text{R}^{2}$C, to address UQ in RAR models. $\text{R}^{2}$C is a consistency-based approach that performs UQ in two main stages, as illustrated in Figure 1: (i) generating the most-likely response, and (ii) sampling multiple generations. The core idea of $\text{R}^{2}$C is to perturb the reasoning paths of these multiple generations through a set of perturbation actions, denoted as $A^{*}$. In MDP terms, $\text{R}^{2}$C temporarily replaces the main action set $A$ with $A^{*}$ for a single state, allowing the RAR model to interleave its generation process and explore new reasoning trajectories, queries, and documents. We define three perturbation actions employed in $\text{R}^{2}$C: (i) Query Paraphrasing, $a_{qp}$, (ii) Critical Rethinking, $a_{cr}$, and (iii) Answer Validation, $a_{av}$. In the following sections, we first formally describe how $\text{R}^{2}$C perturbs the generation path and then describe perturbation actions.
@@ 4.1. $\text{R}^{2}$C: Retrieval-Augmented Reasoning Consistency
$\text{R}^{2}$C models UQ for RAR as an MDP, in which multiple response generations are produced by temporarily replacing the action set $A$ with an alternative set $A^{*}$ at a randomly selected state $s_{t}$. First, the most-likely generation is produced as an MDP iteratively:
@@
The final sampled response $r^{b}$ is obtained from the last state $s_{N^{b}}$ and added to $R_{G}$. After $B$ iterations, we obtain the response set $R_{G}=\{r^{b}\}_{b=1}^{B}$, on which the majority voting function (Eq. (1)) and the uncertainty scoring function (Eq. (2)) are applied to compute the final uncertainty score. Algorithm 1 provides a detailed description of the entire process in $\text{R}^{2}$C.
@@ 4.2. Perturbation Actions
A1: Query Paraphrasing (QP). Constructing effective search queries is critical for retrieving relevant documents; however, LLMs are not inherently optimized for this purpose (Jiang et al., 2025b; Ma et al., 2023; Li et al., 2025b). The QP action ($a_{qp}$) is introduced as a query optimization mechanism that enables the system to explore alternative semantic formulations of the original query. Precisely, when action $a_{qp}$ is applied, the think $\tau_{t}$ of the state $s_{t}$ is preserved and only the query changes. Formally, the LLM $\pi_{\theta}$ takes action $a_{qp}$ at state $s_{t}$, transitioning to state $s_{t+1}$ with the same think $\tau_{t}$ and a new query $q_{t+1}$:
@@
Conceptually, QP perturbation tests whether the reasoning path is so fragile that paraphrasing the search query can alter its direction and lead to the retrieval of different documents. The QP action is implemented by prompting the LLM with a paraphrasing instruction, as shown in Figure 7 in Appendix A.
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025a) ReSearch (Chen et al., 2025c) Search-R1 (Jin et al., 2025) Avg.
@@
A2: Critical Rethinking (CR). RAR models often suffer from the problem of self-criticism, where they fail to recognize that previously retrieved information is noisy or irrelevant (Jin et al., 2025; Jiang et al., 2025b; Asai et al., 2024; Li et al., 2025b). Consequently, they continue to build their reasoning on top of earlier steps, even when those steps are uninformative and lack relevant content. This issue becomes particularly severe when it occurs in the early stages of the reasoning path.
@@
Conceptually, if the reasoning path so far has been incorrect, this action enables the system to adjust toward a more reliable trajectory. If the path has been correct, CR strengthens its validity, thereby increasing confidence in the final outcome. Figure 8 in Appendix A illustrates the instruction prompt used for this action.
A3: Answer Validation (AV). (He et al., 2024; Jiang et al., 2025a; Salemi et al., 2025) RAR models face difficulties with validating their final response, detecting whether the generated response meets certain criteria of the query (Wang et al., 2024; Salemi et al., 2025; He et al., 2024; Jiang et al., 2025a). One challenge is that the response is built upon a reasoning path that integrates both documents and the intermediate reasoning trajectory, which often leads the LLM to exhibit excessive confidence in its output. Another challenge is that different tasks and response types involve specific validation criteria, but RAR models are generally unaware of these requirements.
We introduce the AV action ($a_{av}$) to validate the final response by prompting the LLM to reconsider its generation once a response has been produced, based on predefined criteria. Specifically, the LLM first generates a query-aware reasoning path summary (Shi et al., 2025; Li et al., 2025a; Achkar et al., 2025), and then evaluates the final response using two criteria: (i) Groundedness: is the response supported by the retrieved documents? and (ii) Correctness: does the response appropriately and sufficiently address the query, given the available evidence? Formally, let $D=\{D_{1},\dots,D_{N-1}\}$ be the set of documents retrieved at states $s_{1},s_{2},...s_{N-1}]$. A model $\mathcal{M}$ generates the summary $\hat{S}$ of these documents: $\hat{S}=\mathcal{M}(x,D).$ The state $s_{N}$ is then updated with this summary, denoted as $\hat{s}_{N}$. With the updated state $\hat{s}_{N}$, the LLM $\pi_{\theta}$ is then instructed to generate a new state $s_{N+1}$:
@@
In principle, if the final response $r$ is validated as correct, the system outputs it directly in state $s_{N+1}$. Otherwise, if the validation indicates that the answer is incorrect or incomplete, the system begins a new reasoning path starting from $\langle\tau_{N+1},q_{N+1}\rangle$. The prompt instructions for the path summary model and the action are presented in Figures 9 and 10 in Appendix A, respectively.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Figure 1. $\text{R}^{2}$C overview. Given a user query, the agent (LLM) first generates the most-likely reasoning path leading to the most-likely response (left, yellow). To estimate uncertainty, $\text{R}^{2}$C creates multiple perturbed generations by randomly altering states in the reasoning path (middle, gray). The confidence score is then derived via majority voting. $\text{R}^{2}$C significantly outperforms established UQ methods and achieves significant improvements on two downstream tasks: abstention and model selection.
Retrieval-augmented generation (RAG) is widely used for knowledge-intensive tasks, but remains limited in addressing complex multi-step reasoning (Jin et al., 2025; Soudani et al., 2024b; Salemi and Zamani, 2024; Soudani et al., 2024c). Recent work has explored combining RAG with reasoning, where LLMs are prompted or trained to use search engines as tools during their reasoning process; a paradigm referred to as retrieval-augmented reasoning (RAR) (Sun et al., 2025; Tran et al., 2025; Islam et al., 2024; Jin et al., 2025). However, RAR models are still prone to producing incorrect responses, due to issues such as retrieving irrelevant documents in early steps, misinterpreting retrieved content, or misusing internal knowledge. Therefore, ensuring the trustworthiness of RAR outputs has become a critical challenge.
Uncertainty quantification (UQ) is a widely studied task in machine learning, aimed at assessing the reliability of model outputs by measuring the degree of uncertainty (or lack of confidence)  [^1] The terms uncertainty and confidence are used interchangeably in existing literature. a model has in its predictions (Malinin and Gales, 2021; Kadavath and others, 2022; Kuhn et al., 2023; Hou et al., 2024; Ye et al., 2024; Harrison et al., 2024). Recent methods of estimating the uncertainty of LLM outputs are designed for settings where the input consists solely of a query, meaning the LLM itself is the only source of uncertainty (Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024). The limited work on UQ for RAG (Soudani et al., 2025; Perez-Beltrachini and Lapata, 2025) incorporates the document–response relationship into the uncertainty score, but these methods are only applicable for simple RAG settings, where documents are retrieved once and inserted into the input prompt for generation. As a result, existing UQ approaches are suboptimal for RAR.
@@
In this paper, we propose Retrieval-Augmented Reasoning Consistency ($\text{R}^{2}$C), a novel UQ method that, unlike previous approaches, accounts for multiple sources of uncertainty in RAR. The central idea of $\text{R}^{2}$C is to allow the model to explore diverse reasoning paths, queries, and documents and then measure the consistency of the resulting final answers. This is achieved by modeling RAR as a Markov Decision Process (MDP) and perturbing this process in a controlled way through a set of perturbation actions across various states. Three perturbation actions are designed to influence query generation, document retrieval, and LLMs’ thinking process. These perturbations enable models to arrive at diverse final responses for uncertain generations. The confidence score is then obtained by measuring consistency of the generated answers using majority voting; see Figure 1.
@@
We conduct our experiments across multiple datasets and five RAR models. Our experiments show that $\text{R}^{2}$C significantly outperforms existing LLM-specific UQ methods, achieving on average more than a 5% improvement in AUROC compared to the state-of-the-art UQ methods. We further extrinsically evaluate $\text{R}^{2}$C on two downstream tasks: (i) Abstention: [^2] Also referred to as selective prediction in the literature (Duan et al., 2025; Xin et al., 2021). the task of generating ‘I don’t know’ when the model is uncertain about its output (Feng et al., 2024; Madhusudhan et al., 2025), and (ii) Model Selection: [^3] Also referred to as selection-based model ensemble in the literature (Jiang et al., 2023a). the task of selecting a final answer from a pool of candidates generated by multiple systems (Chen et al., 2025c; Guha et al., 2024). Our experimental results indicate that $\text{R}^{2}$C delivers statistically significant gains over existing approaches: in Abstention, it achieves roughly 5% improvements on both F1Abstain and AccAbstain; in Model Selection, it increases exact match by about 7% relative to single RAR models and by about 3% compared to selection model baselines.
@@
Given the strong performance of $\text{R}^{2}$C in both direct evaluation and extrinsic evaluation on downstream tasks, we investigate the factors that contributes to its effectiveness. We show that $\text{R}^{2}$C retrieves on average 25 unique documents for each score, compared to 16 documents retrieved by other UQ methods. It also achieves a query diversity of $0.35$ compared to $0.30$ other methods, measured by the inverse of the average pairwise cosine similarity between queries (Cox et al., 2021; Zhang et al., 2025). This diversity in queries and documents demonstrates that $\text{R}^{2}$C generates diverse, yet relevant reasoning paths through our controlled perturbation mechanism. As a result, this enables the method to achieve confidence scores comparable to baseline approaches while requiring only about 3 generations on average, 2.5 times fewer token generations than the 10 used by the baselines. This highlights that $\text{R}^{2}$C is not only the most effective method of its kind but also a relatively more efficient UQ approach. To summarize, the main contributions of this paper are:
@@ 2. Related Work
Retrieval-Augmented (Reasoning) Models. RAG is a framework that combines the strengths of retrieval models and generative models (Cuconasu et al., 2024). Broadly, RAG can be implemented in different ways. In the retrieve-then-generate paradigm, relevant documents are first retrieved based on the user’s input and then incorporated into the model’s prompt (Mallen et al., 2023; Soudani et al., 2024a). In contrast, Active RAG allows retrieval to occur throughout the generation process, either in fixed intervals or dynamically, whenever additional information is needed (Trivedi et al., 2023; Jiang et al., 2023b; Su et al., 2024). Retrieval-Augmented Reasoning (RAR) is a recent extension of RAG that integrates retrieval with reasoning, aiming to improve the interaction between LLMs and retrievers (Sun et al., 2025; Tran et al., 2025). For example, SelfAsk (Press et al., 2023) decomposes complex questions into follow-up queries and intermediate answers, while ReAct (Yao et al., 2023) defines a set of actions, such as search, look up, and finish, to structure interactions with external resources. More recent and effective models such as ReSearch (Chen et al., 2025a) and Search-R1 (Jin et al., 2025) are explicitly trained to seamlessly integrate external resources into reasoning. Despite their effectiveness on complex queries, RAR models remain prone to errors, such as retrieving irrelevant documents, misinterpreting content, or misusing internal knowledge.
Uncertainty Quantification (UQ) and Confidence Estimation (CE). UQ and CE are closely related but conceptually distinct. In traditional machine learning, uncertainty is defined as a property of the model’s predictive distribution given a specific input (Lin et al., 2024; Liu et al., 2025). Confidence, by contrast, reflects the model’s belief in the correctness of a particular prediction for the given input. In other words, CE is defined as the task of quantifying how certain a model is about a specific generated response (Liu et al., 2025), while UQ is a task of capturing the degree of variability or unpredictability in the model’s outputs (irrespective of a specific response) (Kendall and Gal, 2017; Gawlikowski et al., 2023). Despite this distinction, existing work sometimes use uncertainty to refer to estimated confidence (Kadavath and others, 2022; Liu et al., 2025). In this paper, we compute confidence scores for LLM generated responses. Nevertheless, following common practice in the literature (Kadavath and others, 2022; Malinin and Gales, 2021; Wang et al., 2024b; Ren et al., 2022; Fadeeva et al., 2024; Yin et al., 2024; Zhang and Zhang, 2025; Becker and Soatto, 2024; Liu et al., 2024), we refer to this confidence score as an uncertainty measure and compare it against other UQ methods.
@@
Broadly, existing UQ methods can be divided into two categories: white-box approaches, which leverage token-level probabilities and entropy (Kadavath and others, 2022; Kuhn et al., 2023; Bakman et al., 2024; Yaldiz et al., 2025a; Duan et al., 2024; Soudani, 2025), and black-box approaches, which rely only on the final textual outputs (Lin et al., 2024; Tian et al., 2023). Most UQ methods focus on question answering and view the LLM as the only source of uncertainty, but in RAG, the retriever also contributes its own uncertainty, making this assumption incomplete. Limited research has explored UQ for RAG by modeling the document-response link, either via axioms (Soudani et al., 2025) or utility models (Perez-Beltrachini and Lapata, 2025). Some path-based approaches focus on assessing the consistency of reasoning paths in reasoning tasks (Wang et al., 2023; Qi et al., 2025; Li et al., 2025a). However, these methods do not extend naturally to RAR that involves repeated retrieval during reasoning. Recent work has studied uncertainty propagation in multi-step decision-making by combining uncertainties from intermediate steps. SAUP (Zhao et al., 2025) learns aggregation weights to merge per-step uncertainties, but it relies on ground-truth labels from the test domain. In contrast, we propose $\text{R}^{2}$C, a method that accumulates uncertainty over the entire reasoning path, while considering different sources of uncertainty including the retriever and the generator.
@@ 3. Preliminaries
Consistency-based Uncertainty Quantification. The core idea of consistency-based methods (Xiong et al., 2023; Hou et al., 2024) is to generate multiple responses for a given query by varying either the input prompt or the temperature parameter used in stochastic decoding (Korikov et al., 2025). The idea build upon the hypothesis of self-consistency theory, where correct reasoning processes, even if they are diverse, tend to have greater agreement in their final answer (Wang et al., 2023). The pairwise similarity among these responses is then computed and aggregated into a single confidence score (Lin et al., 2024; Bakman et al., 2025). Formally, consider a model $\pi$ parameterized with $\theta$, generates the most-likely response $r$ by setting the sampling temperature to less than one. Then, $B$ additional responses $R_{G}=\{r^{b}\}_{b=1}^{B}$ are sampled using various sampling strategies such as increasing the temperature (Wang et al., 2023), changing the input (Korikov et al., 2025), or altering the reasoning path (Qi et al., 2025; Li et al., 2025a). A transformation function $\mathcal{\phi}$ is then applied to convert $R_{G}$ into a confidence score. Our method builds on the reasoning path perturbation approach and employs Majority Voting (Wang et al., 2023) as the transformation function, where confidence is the degree of consistency, measured by the proportion of sampled responses that match the most-likely response:
@@ 4. Methodology
We propose Retrieval-Augmented Reasoning Consistency, $\text{R}^{2}$C, to address UQ in RAR models. $\text{R}^{2}$C is a consistency-based approach that performs UQ in two main stages, as illustrated in Figure 1: (i) generating the most-likely response, and (ii) sampling multiple generations. The core idea of $\text{R}^{2}$C is to perturb the reasoning paths of these multiple generations through a set of perturbation actions, denoted as $A^{*}$. In MDP terms, $\text{R}^{2}$C temporarily replaces the main action set $A$ with $A^{*}$ for a single state, allowing the RAR model to interleave its generation process and explore new reasoning trajectories, queries, and documents. We define three perturbation actions employed in $\text{R}^{2}$C: (i) Query Paraphrasing, $a_{qp}$, (ii) Critical Rethinking, $a_{cr}$, and (iii) Answer Validation, $a_{av}$. By employing a variety of actions to perturb the reasoning processes of LLMs, $\text{R}^{2}$C captures both epistemic (model) uncertainty and aleatoric (data) uncertainty. This perspective is supported by Liu et al. (2025), who argue that reasoning uncertainty encompasses both epistemic and aleatoric components. In this work, we do not seek to disentangle these two sources of uncertainty; rather, we measure their combined effect as total uncertainty.
We note at the outset that $\text{R}^{2}$C builds on the following definition for perturbation actions:
Definition. A perturbation action changes the state of a system without omitting existing information or adding extra knowledge to the system.
This aligns our method with the self-consistency theory and ensures that any divergence between the most-likely response and the generated response after applying a perturbation action is attributable to the variability of stochastic response generation, rather than to changes in the system’s knowledge. In the following sections, we first formally describe how $\text{R}^{2}$C perturbs the generation path and then describe perturbation actions.
@@ 4.1. $\text{R}^{2}$C: Retrieval-Augmented Reasoning Consistency
$\text{R}^{2}$C models uncertainty for RAR as an MDP, in which multiple response generations are produced by temporarily replacing the action set $A$ with an alternative set $A^{*}$ at a randomly selected state $s_{t}$. First, the most-likely generation is produced as an MDP iteratively:
@@
The final sampled response $r^{b}$ is obtained from the last state $s_{N^{b}}$ and added to $R_{G}$. After $B$ iterations, we obtain the response set $R_{G}=\{r^{b}\}_{b=1}^{B}$, on which the confidence score function (Eq. (1)) is applied to compute the final confidence score. Algorithm 1 provides a detailed description of the entire process in $\text{R}^{2}$C.
@@ 4.2. Perturbation Actions
A1: Query Paraphrasing (QP). Constructing effective search queries is critical for retrieving relevant documents; however, LLMs are not inherently optimized for this purpose (Jiang et al., 2025b; Ma et al., 2023; Li et al., 2025a). The QP action ($a_{qp}$) is introduced as a query optimization mechanism that enables the system to explore alternative semantic formulations of the original query. Precisely, when action $a_{qp}$ is applied, the think $\tau_{t}$ of the state $s_{t}$ is preserved and only the query changes. Formally, the LLM $\pi_{\theta}$ takes action $a_{qp}$ at state $s_{t}$, transitioning to state $s_{t+1}$ with the same think $\tau_{t}$ and a new query $q_{t+1}$:
@@
Conceptually, QP perturbation tests whether the reasoning path is so fragile that paraphrasing the search query can alter its direction and lead to the retrieval of different documents. The QP action is implemented by prompting the LLM with a paraphrasing instruction.
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
@@
A2: Critical Rethinking (CR). RAR models often suffer from the problem of self-criticism, where they fail to recognize that previously retrieved information is noisy or irrelevant (Jin et al., 2025; Jiang et al., 2025b; Asai et al., 2024; Li et al., 2025a). Consequently, they continue to build their reasoning on top of earlier steps, even when those steps are uninformative and lack relevant content. This issue becomes particularly severe when it occurs in the early stages of the reasoning path.
@@
Conceptually, if the reasoning path so far has been incorrect, this action enables the system to adjust toward a more reliable trajectory. If the path has been correct, CR strengthens its validity, thereby increasing confidence in the final outcome.
A3: Answer Validation (AV). RAR models face difficulties with validating their final response, detecting whether the generated response meets certain criteria of the query (Wang et al., 2024a; Salemi et al., 2025; He et al., 2024; Jiang et al., 2025a). One challenge is that the response is built upon a reasoning path that integrates both documents and the intermediate reasoning trajectory, which often leads the LLM to exhibit excessive confidence in its output. Another challenge is that different tasks and response types involve specific validation criteria, but RAR models are generally unaware of these requirements.
We introduce the AV action ($a_{av}$) to validate the final response by prompting the LLM to reconsider its generation once a response has been produced, based on predefined criteria. Specifically, the LLM first generates a query-aware reasoning path summary (Shi et al., 2025; Li et al., 2025b; Achkar et al., 2025), and then evaluates the final response using two criteria: (i) Groundedness: is the response supported by the retrieved documents? and (ii) Correctness: does the response appropriately and sufficiently address the query, given the available evidence? Formally, let $D=\{D_{1},\dots,D_{N-1}\}$ be the set of documents retrieved at states $s_{1},s_{2},...s_{N-1}]$. A model $\mathcal{M}$ generates the summary $\hat{S}$ of these documents: $\hat{S}=\mathcal{M}(x,D).$ The state $s_{N}$ is then updated with this summary, denoted as $\hat{s}_{N}$. With the updated state $\hat{s}_{N}$, the LLM $\pi_{\theta}$ is then instructed to generate a new state $s_{N+1}$:
@@
In principle, if the final response $r$ is validated as correct, the system outputs it directly in state $s_{N+1}$. Otherwise, if the validation indicates that the answer is incorrect or incomplete, the system begins a new reasoning path starting from $\langle\tau_{N+1},q_{N+1}\rangle$.